OpenStudy (anonymous):

Urgent please help asap!! Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -1/3x - 6 y = -1/3x - 14 y = 3x - 6 y = 3x - 14

OpenStudy (igreen):

First rearrange the equation: 3x - y = -3 Subtract 3x to both sides: -y = -3x - 3 Multiply -1 to all terms: y = 3x + 3 Okay, so now our equation is in the form of y = mx + b, where m = slope and b = y-intercept. y = mx + b y = 3x + 3 Therefore, the slope of this equation is 3, and the y-intercept is also 3. That already eliminates the first two choices.

OpenStudy (anonymous):

https://www.desmos.com/calculator This shall help for future questions.

OpenStudy (igreen):

Now to find the answer, plug in (4, -2) into the last two equations and see which one fits: y = 3x - 6 -2 = 3(4) - 6 Multiply: -2 = 12 - 6 Subtract: -2 = 6 (4, -2) does NOT fit in the equation y = 3x - 6. y = 3x - 14 -2 = 3(4) - 14 Multiply: -2 = 12 - 14 Can you subtract that?

OpenStudy (igreen):

@xxbeautifultragedyxx

OpenStudy (anonymous):

wait didn't you just subtract that? -2=12-14

OpenStudy (igreen):

Yes. Subtract 12 - 14, what do you get?

OpenStudy (anonymous):

-2

OpenStudy (igreen):

Yes! So you get -2 = -2, and that's true. And since (4, -2) fits in the equation y = 3x - 14, it is the answer.

OpenStudy (anonymous):

Thank you! could you help me with 1 more?

OpenStudy (igreen):

Sure thing, just tag me.